# Evaluation for Regression Analyses on Evolving Data Streams

### v1

@@ Abstract.
The paper explores the challenges of regression analysis in evolving data streams, an area that remains relatively underexplored compared to classification. We propose a standardized evaluation process for regression and prediction interval tasks in streaming contexts. Additionally, we introduce an innovative drift simulation strategy capable of synthesizing various drift types, including the less-studied incremental drift. Comprehensive experiments with state-of-the-art methods, conducted under the proposed process, validate the effectiveness and robustness of our approach.
@@
A demonstration of the simulation of incremental drifts with a synthetic dataset that includes one feature column and one target column.
@@ 1. Introduction
Machine learning on streaming data has garnered significant interest due to its applicability in dynamic and evolving environments (Gunasekara et al., 2023; Read and Zliobaite, 2023). However, while extensive research has focused on classification tasks in stream learning, regression tasks remain underexplored (Choudhary et al., 2021), one of the reason is the lack of dedicated data resources (Sun et al., 2024b). The challenges are further compounded by the difficulty in defining and identifying the drift in the real-world data sequences (Wares et al., 2019). A real-world application of data stream regression includes real-time energy pricing adjustments based on evolving market conditions (Sun et al., 2024a). Additionally, Prediction Intervals (PI) for streaming data have demonstrated significant effectiveness in quantifying regression uncertainty (Sun et al., 2024a).
@@ 2.3. Real-World Datasets
Abalone (Nash et al., 1995) is a well-known dataset from the UCI Machine Learning Repository. It contains measurements of physical attributes of abalones, such as length, diameter, weight, and shell dimensions, aimed at predicting the age of the abalones. The Bike Sharing dataset (Fanaee-T and Gama, 2014) is a commonly used regression dataset that provides data on bike rental demand, integrating environmental and seasonal settings with temporal features. The House8L dataset (Torgo, 1990), derived from the 1990 U.S. Census, is designed for regression tasks aimed at predicting the median house prices in various regions. The Superconductivity dataset, presented by Hamidieh in (Hamidieh, 2018), contains data for predicting the critical temperature of superconductors based on material properties. Real-time energy price datasets from (Sun et al., 2024a) capture dynamic electricity pricing and consumption patterns in New Zealand. These datasets include real-time pricing, historical trends, and contextual variables like weather and time-of-use, emphasizing the role of real-time analysis in managing energy demand and supply for sustainability and cost-effectiveness. Consistently updated pricing data ensures relevance for research and application.
@@ 3. Problem Definition
Regressive analysis in stream learning focuses on building and maintaining models capable of predicting continuous target variables from evolving data streams. Unlike traditional batch learning, stream learning must accommodate challenges such as real-time processing, limited memory, and concept drifts. Regression tasks aim to provide accurate point predictions for the target variable, which are crucial for decision-making in dynamic environments. Beyond point predictions, stream learning also emphasizes uncertainty quantification through Prediction Intervals (PIs), which define a range within which the true target value is expected to lie with a given confidence level.
3.1. Regression
Regression tasks in data stream learning involve predicting a continuous target variable from an incoming, potentially infinite stream of data. These tasks pose unique challenges, including the need to handle evolving data distributions, limited memory, and real-time processing constraints.
3.1.1. Evaluation Methods
To assess the performance of regression models in a data stream setting, two primary evaluation approaches are commonly used:
@@ 3.2. Prediction Interval
Prediction Intervals (PIs) are critical in stream learning scenarios to quantify the uncertainty associated with regression predictions. A PI provides a range within which the true value of the target variable is expected to lie with a specified confidence level. This is particularly important in dynamic data streams where uncertainty may evolve over time.
@@ 3.2.1. Evaluation Metrics
Two key metrics are used to evaluate the quality of PIs in stream learning:
@@ 4. Related Works
Regressive analyses have always been overlooked in Machine Learning field, not to mention the contempt for the streaming scenario. Novel research on regression and prediction intervals is limited in the recent literature. We choose some new and commonly used algorithms as demonstration tools in this work.
@@ 6.1. Algorithms and Parametrization
The following configurations were used to ensure fair comparison and reproducibility: (1) Sliding Window KNN: $k=10$, window size = 1000. (2) FIMT-DD: Grace period = 200, split confidence = 0.01. (3) ARF-Regand SOKNL: Ensemble size = 30. These settings align with common literature practices and ensure effective performance across datasets. For Prediction Interval experiments, both MVE and AdaPI (Section 4) used a 95% confidence level. AdaPI’s lower limit was set to 0.01, while KNN and SOKNL served as base models with the same configurations. Prequential executions used a window size of 1000. All participants can be located at MOA (Bifet et al., 2010) and CapyMOA, both of which are open-sourced stream learning platforms.
@@ 6.2. Regression
Figure 3 compares four regression algorithms – KNN, FIMT-DD, ARF-Reg, and SOKNL – across 18 datasets using their cumulative Adjusted $\mathcal{R}^{2}$ ($\mathcal{R}^{2}_{adj}$). The x-axis represents the datasets, separated into groups by the vertical pink dashed lines. Each group has the same source of original data, i.e., the first group is derived from Abalone dataset. The standard deviations from ten executions are significantly small, thus omitted in the figure. Because different datasets yield different scale of RMSE ($\sigma_{e}$) results, it is impossible to illustrate all results within a single figure. Therefore, we avoid the illustration for RMSE in the main contents of this paper. Full $\mathcal{R}^{2}_{adj}$ and $\sigma_{e}$ results can be located in the Appendix (Table 3).
@@
Figure 5 is a prequential $\mathcal{R}^{2}_{adj}$ results demonstration for the NZEP dataset group. A key observation is that for the abrupt and gradual (Figure 5(a) and 5(b)), only the last concept cause an significant difference. In our case, the fourth concept represents the energy price in Dunedin, which is the only city from the South Island of New Zealand. This indicates a potential difference in the pricing strategy between North (AKL, HAM, and WEL) and South Island (DUN). In terms of the incremental ones (Figure 5(c) – 5(f)), they all illustrate different degrees of unstability due to the incremental simulated with the most correlated feature to the target values.
@@ 6.3. Prediction Interval
Figure 6 and 7 present the cumulative coverage ($\mathcal{C}$) and normalized mean prediction interval width ($\mathcal{W}_{norm}$) results for different datasets, comparing the performance of MVE and AdaPI methods with kNN and SOKNL models. A tabular result summarization of the cumulative $\mathcal{C}$ and $\mathcal{W}_{norm}$ is situated at Table 5 in the Appendix.
Figure 6. Coverage ($\mathcal{C}$) for Two Prediction Interval Algorithms with Two Base Regressors on 18 Datasets
@@
After encountering the second drift, the KNN struggles to represent the new concept, leading to a loss of accuracy. In response, AdaPI attempts to recover Coverage by widening the generated intervals to capture more ground truths. Consequently, the NMPIW for AdaPI becomes noticeably larger than that of MVE, as depicted in Figure 8(b). The outcome is reflected in Figure 8(a), where the Coverage for AdaPI regains the 95% at a faster pace.
@@
If we take a more focus look into the streaming progress, as demonstrated in Figure 9, which is a clip of the entire stream of the generated PI areas, it can be spotted that AdaPI (red shadowed region) is larger than the MVE (blue shadowed region) and covers more target values (the orange “+” symbols). This behaviour makes sure that the coverage recovers to the desired values more rapidly.
Figure 9. A Clip of the Prediction Interval generated by MVE and AdaPI (from 120,000 to 120,200)

### v2

# Evaluation for Regressive Analyses on Evolving Data Streams

@@ Abstract.
This paper explores the challenges of regression analysis in evolving data streams, an area that remains relatively underexplored compared to classification. We propose a standardized evaluation process for regression and prediction interval tasks in streaming contexts. Additionally, we introduce an innovative drift simulation strategy capable of synthesizing various drift types, including the less-studied incremental drift. Comprehensive experiments with state-of-the-art methods, conducted under the proposed process, validate the effectiveness and robustness of our approach.
@@
A demonstration of the simulation of incremental drifts with the bike dataset that includes one feature column and one target column.
@@ 1. Introduction
Machine learning on streaming data has garnered significant interest due to its applicability in dynamic and evolving environments (Gunasekara et al., 2023; Read and Zliobaite, 2023). However, while extensive research has focused on classification tasks in stream learning, regression tasks remain underexplored (Choudhary et al., 2021), with one of the reasons being a lack of dedicated data resources (Sun et al., 2024b). The challenges are further compounded by the difficulty in defining and identifying drift in the real-world data sequences (Wares et al., 2019). A good example for real-world applications of data stream regression is real-time energy pricing adjustments based on evolving market conditions (Sun et al., 2024a). In addition to a point prediction, regression also generally needs uncertainty quantification, which can be provided by Prediction Intervals (PI) for streaming data (Sun et al., 2024a).
@@ 2.3. Real-World Datasets
Abalone (Nash et al., 1995) is a well-known dataset from the UCI Machine Learning Repository. It contains measurements of physical attributes of abalones, such as length, diameter, weight, and shell dimensions, aimed at predicting the age of the abalones. The Bike Sharing dataset (Fanaee-T and Gama, 2014) is a commonly used regression dataset that provides data on bike rental demand, integrating environmental and seasonal settings with temporal features. The House8L dataset (Torgo, 1990), derived from the 1990 U.S. Census, is designed for regression tasks aimed at predicting the median house prices in various regions. The Superconductivity dataset, presented by Hamidieh in (Hamidieh, 2018), contains data for predicting the critical temperature of superconductors based on material properties. Real-time energy price datasets from (Sun et al., 2024a) capture dynamic electricity pricing and consumption patterns in New Zealand. These datasets include real-time pricing, historical trends, and contextual variables, emphasizing the role of real-time analysis in managing energy demand and supply for sustainability and cost-effectiveness. Consistently updated pricing data ensures relevance for research and application.
@@ 3. Problem Definition
Regressive analysis in stream learning focuses on building and maintaining models capable of predicting continuous target variables from evolving data streams. Unlike traditional batch learning, stream learning must accommodate challenges such as real-time processing, limited memory, and concept drifts. Regression tasks aim to provide accurate point predictions for the target variable, which are crucial for decision-making in dynamic environments. Beyond point predictions, stream learning also emphasizes uncertainty quantification through Prediction Interval (PI), which define a range within which the true target value is expected to lie with a given confidence level.
3.1. Evaluation Protocols
Evaluation under streaming setting is different from conventional machine learning. In addition to the overall performance, the trends and fluctuations during the processing of a stream are also vital (Gunasekara, 2023; Bifet et al., 2010). With the drift detection and adaptation abilities, the streaming algorithms ought to recover the performance from the presence of concept drifts, which will not be reflected in the overall evaluations. Thus, in stream learning, researchers usually present two evaluation protocols, Cumulative and Prequential (Gomes et al., 2025).
@@ 3.1. Evaluation Protocols
3.2. Regression
Literature has proposed plenty of evaluation metrics for regression tasks, most of which are either error-based or correlation-based (Chicco et al., 2021). We recommend one of each in this work.
@@ 3.3. Prediction Interval
As presented in (Sun et al., 2024b; Shafer and Vovk, 2008), evaluating prediction intervals is a challenging task. Extremely wide intervals may cover more ground truths but significantly reduce informativeness. Balancing this trade-off requires optimizing two key metrics.
@@ 3.3.1. Evaluation Metrics
In the literature ((Hadjicharalambous et al., 2020; Pang et al., 2018; Ozek et al., 2024; Zhong and Xu, 2021)), two key metrics are commonly used to evaluate the quality of PIs in stream learning:
@@ 4. Related Works
Regression analyses have always been overlooked in Machine Learning field, not to mention the contempt for the streaming scenario. Novel research on regression and prediction intervals is limited in the recent literature. We choose some new and commonly used algorithms as demonstration tools in this work.
@@ 6.1. Algorithms and Parametrization
The following configurations are provided to ensure and reproducibility: (1) Sliding Window KNN: $k=10$, window size = 1000. (2) FIMT-DD: Grace period = 200, split confidence = 0.01. (3) ARF-Regand SOKNL: Ensemble size = 30. These settings align with common literature practices and ensure effective performance across datasets. For Prediction Interval experiments, both MVE and AdaPI (Section 4) used a 95% confidence level. AdaPI’s lower limit was set to 0.01, while KNN and SOKNL served as base models with the same configurations. Prequential executions used a window size of 1000. All participants can be easily accessed from open-sourced stream learning platforms, such as MOA (Bifet et al., 2010) and CapyMOA (Gomes et al., 2025).
@@ 6.2. Regression
Figure 3 compares four regression algorithms – KNN, FIMT-DD, ARF-Reg, and SOKNL – across 18 datasets using their cumulative Adjusted $\mathcal{R}^{2}$ ($\mathcal{R}^{2}_{adj}$). The x-axis represents the datasets, separated into groups by the vertical pink dashed lines. Each group has the same source of original data, e.g., the first group is derived from Abalone dataset. The standard deviations from ten executions are significantly small, thus omitted in the figure. Because different datasets yield different scale of RMSE ($\sigma_{e}$) results, it is unfeasible to illustrate all $\sigma_{e}$ results within a single figure without rescaling. Therefore, we avoid the illustration for RMSE in the main contents of this paper. Full $\mathcal{R}^{2}_{adj}$ and $\sigma_{e}$ results can be located in Table 3 in Appendix A.1.
@@
Figure 5 is a prequential $\mathcal{R}^{2}_{adj}$ result demonstration for the NZEP dataset group. A key observation is that for the abrupt and gradual (Figure 5(a) and 5(b)), only the last concept cause an significant difference. In our case, the fourth concept represents the energy price in Dunedin, which is the only city from the South Island of New Zealand. This indicates a potential difference in the pricing strategy between North (AKL, HAM, and WEL) and South Island (DUN). In terms of the incremental ones (Figure 5(c) – 5(f)), in all of which different degrees of unstability can be observed. The reason is likely to be the effectiveness of the incremental drifts simulated with the most correlated feature to the target values.
@@ 6.3. Prediction Interval
Figure 6 and 7 present the cumulative coverage ($\mathcal{C}$) and NMPIW ($\mathcal{W}_{norm}$) results for different datasets, comparing the performance of MVE and AdaPI methods with KNN and SOKNL models. A tabular result summarization of the cumulative $\mathcal{C}$ and $\mathcal{W}_{norm}$ is situated at Table 5 in the Appendix A.2.
Figure 6. Coverage ($\mathcal{C}$) for Two Prediction Interval Algorithms with Two Base Regressors on 18 Datasets. The red dashed line highlights the confidence level, which the PI methods aim to be closer to
@@
After encountering the second drift, the KNN struggles to represent the new concept, leading to an increase in errors. In response, AdaPI attempts to recover Coverage by widening the generated intervals to capture more ground truths. Consequently, the NMPIW for AdaPI becomes noticeably larger than that of MVE, as depicted in Figure 8(b). The outcome is reflected in Figure 8(a), where the Coverage for AdaPI regains the 95% at a faster pace.
Figure 9. A Clip of the Prediction Interval from MVE and AdaPI after the Second Drift (Instance #120,000 to #120,200)
If we take a more focus look into the streaming progress, as demonstrated in Figure 9, which is a clip of the entire stream of the generated PI areas, it can be spotted that AdaPI (red shadowed region) is larger than the MVE (blue shadowed region) and covers more target values (the additional truths covered by AdaPI are marked by orange “+” symbols). This behaviour makes sure that the coverage recovers to the desired confidence level more rapidly.